OpenStudy (ajprincess):

Please help:) if a and b are two integers with g.c.d(a,b)=1, then show that g.c.d(a-b, a^2+ab+b^2)=1 or 3

OpenStudy (ajprincess):

Let d=g.c.d(a-b,a^2+ab,b^2). We want to show that d=1 or 3. Since d is a divisor of a-b it is also a divisor of (a-b)^2 and -(a-b)^2. Since d is a divisor of -(a-b)^2 and a^2+ab+b^2 it is a divisor of their sum. Sum: -(a-b)^2+a^2+ab+b^2=3ab Am I right upto this part?

OpenStudy (anonymous):

Hello ajprincess quite right till there..

OpenStudy (ajprincess):

Hello. Thanx @mukushla. Hw will I prove d=1 or 3 after this?

OpenStudy (ajprincess):

=> d|3ab => d|3a and d|b or d|a and d|3b Thus d=1 or 3 I am nt sure if I am right

OpenStudy (anonymous):

guys, lend a hand :) what do u think?

OpenStudy (shubhamsrg):

You know there is a saying, if you don't know much about something, better keep mum and let others wonder whether you know it or not, rather than opening your mouth and letting others know! I'd like to follow that principle! :)

hartnn (hartnn):

i think you need to show that d is divisor of ab also.

hartnn (hartnn):

then since d is divior of ab and 3ab, d=1 or 3

hartnn (hartnn):

do u know how can u prove that di is divisor of ab ?

OpenStudy (ajprincess):

hmm no.

hartnn (hartnn):

d is divisor of (a-b)^2 and a^2+ab+b^2 take difference this time...

hartnn (hartnn):

i was doing that in mind...so just a sec...

hartnn (hartnn):

you'll again get 3ab :P

OpenStudy (ajprincess):

ya it will be -3ab this time

mathslover (mathslover):

wait for a second friends ...

OpenStudy (sirm3d):

suppose d|a, because d|a-b then d|b d|a, d|b, d^2|ab d is the greatest, and d^2|ab then d=1

OpenStudy (anonymous):

u may do it like this \(d|a-b\) and \(d|3ab\) so \(d|3ab-3b(a-b)=3b^2\) using same procedure \(d|a-b\) and \(d|3ab\) so \(d|3ab+3a(a-b)=3a^2\) finally \(d|\gcd(3a^2,3b^2)=3\gcd(a^2,b^2)=3\)

OpenStudy (ajprincess):

Sorry. I dnt get the part hw u got d|3ab-3b(a-b)

mathslover (mathslover):

Given : (a,b) = 1 To prove : \(\large{(a-b),(a^2+b^2+ab)= 1 \space \textbf{or} \space 3}\) proof : (a,b) = 1 , let : a = bq + r , \(\large{0 < r < q }\) and (b,r) = 1 are satisfied a - b = bq + r - b = b(q-1) + r let \(\large{(a-b),(a^2+b^2+ab) \ne 1}\) => \(\large{a-b|a^2 + b^2 + ab}\) --------------1) Also , \(\large{a-b | (a-b)^2}\) => \(\large{a-b|a^2+b^2-2ab}\) ----2) From 1) and 2) : \(\large{a-b|a^2 + b^2 + ab - a^2 -b^2 +2ab \implies a-b |3ab}\) Now , \(\large{a-b \space \textbf{not divide} \space a}\) and \(\large{a-b \space \textbf{not divide} \space b}\) therefore \(\large{a-b | 3}\) Hence \(\large{ (a-b, a^2+b^2+ab) = 3}\) iff it is not equal to 1 now, if \(\large{a-b \space \textbf{not divide} \space 3}\) then : a-b not divide 3ab therefore a-b not divides (a-b)^2 + 3ab implies, a-b not divides (a^2 +b^2 + ab) that is (a-b, a^2+b^2+ab) = 1 proved

OpenStudy (anonymous):

@ajprincess when d is divisor of both a and b then d is also a divisor of every linear combination of a and b like am+bn

mathslover (mathslover):

Is my method correct ?

OpenStudy (ajprincess):

Sorry @mathslover I havnt any idea if ur method is right. i got it @mukushla. so first we need to show d is a divisor of a and b right?

hartnn (hartnn):

no, i think, what he meant was, if d is divisor of A and B, then d is also divisor of A and -3b(B) here A =3ab, B=a-b

mathslover (mathslover):

np , even I came to the above solution after thinking a lot... May be you didn't get it because I posted the solution in a single post and didn't go through step by step and making you understand that by discussing, I was late and so I thought I should post it in a single post now. Sorry but nice work @mukushla

OpenStudy (ajprincess):

Thanx a lot @mathslover, @mukushla nd @hartnn